### v4

# Adversarial Attacks to Latent Representations of Distributed Neural Networks in Split Computing

Abstract
Distributed deep neural networks (DNNs) have been shown to reduce the computational burden of mobile devices and decrease the end-to-end inference latency in edge computing scenarios. While distributed DNNs have been studied, to the best of our knowledge, the resilience of distributed DNNs to adversarial action remains an open problem. In this paper, we fill the existing research gap by rigorously analyzing the robustness of distributed DNNs against adversarial action. We cast this problem in the context of information theory and rigorously proved that (i) the compressed latent dimension improves the robustness but also affect task-oriented performance; and (ii) the deeper splitting point enhances the robustness but also increases the computational burden. These two trade-offs provide a novel perspective to design robust distributed DNN. To test our theoretical findings, we perform extensive experimental analysis by considering 6 different DNN architectures, 6 different approaches for distributed DNN and 10 different adversarial attacks using the ImageNet-1K dataset.
keywords:
Distributed DNN , Adversarial Robustness , Split Computing
journal: Computer Networks
[…]
[label1]organization=Institute for the Wireless Internet of Things at Northeastern University, country=United States
[…]
[label2]organization=Air Force Research Laboratory, country=United States
1 Introduction
Deep neural networks (DNNs) have achieved significant success in various domains such as computer vision [1], natural language processing [2], and wireless communication [3], among many others. However, state-of-the-art DNNs are challenging to deploy on resource-limited mobile devices. While mobile-specific DNNs have been proposed [4], they usually come with a significant loss in accuracy. On the other hand, completely offloading the computation to edge or cloud computers is impractical in mobile scenarios due to the excessive communication overhead corresponding to the transfer of the DNN input from the mobile device to the edge/cloud [5]. A new paradigm called distributed computing – also referred to as split computing in prior art [6] – divides the computation of DNNs across multiple devices – according to the available processing power and networking bandwidth. The key advantage is that optimal load distribution can be achieved while meeting maximum end-to-end latency constraints and also preserving the DNN accuracy [7]. For an excellent survey on distributed/split computing, the reader is referred to the work by [6].
Without loss of generality, we assume that a DNN model is divided into a mobile DNN and a local DNN, respectively executed by the mobile device and an edge/cloud computer. Usually, the DNN architecture is modified by introducing a compression layer ( “bottleneck”) at the end of the mobile DNN [8, 9, 10, 11, 12], which is trained to learn a latent representation that reduces the amount of data being sent to the edge/cloud. The compressed representation is then used by the local DNN to produce the final prediction output (e.g., classification).
Although prior work has proven the advantages of distributing the DNN computation, it is also evident that this approach opens the door to adversarial attacks to intermediate (latent) representations. The distributed nature of the computation exposes the latent representation to adversarial action. Indeed, due to the need for communicating the latent representation across devices over a wireless network, an adversary can easily eavesdrop the latent representation and craft an adversarial sample to compromise the local DNN as shown in Figure 1.
Figure 1: Overview of Adversarial Attacks to Distributed DNNs.
Despite its significance and timeliness, to the best of our knowledge, assessing the robustness of distributed DNNs remains an unexplored problem. We remark that achieving a fundamental understanding of these attacks and evaluating their effectiveness in state-of-the-art DNNs is paramount to design robust distributed DNNs. To this end, we theoretically analyze the robustness of distributed DNNs using Information Bottleneck (IB) theory [13]. Our first key theoretical finding is that the latent robustness is highly depended on the depth of splitting point. With similar levels of information distortion, latent representations in deeper layers exhibit constantly better robustness compare to those in early layers. Our second key finding is that the trade-off between DNN performance and robustness is intrinsically related to the dimension of the latent space. While a smaller latent space may increase robustness by reducing the model variance, it will also introduce bias in the model thus affecting the generalization capability of the DNN model.
We extensively evaluate our theoretical findings by considering 10 adversarial algorithms, i.e., 4 white-box attacks [14, 15, 16, 17] and 6 black-box attacks [18, 19, 20, 21, 22, 23]. We apply these attacks to 6 reference architectures [24, 25] designed with 6 distributed DNN approaches [26, 11, 12, 7, 27, 28]. The experimental results validate our theoretical findings on the examined DNNs and attack algorithms.
Summary of Novel Contributions
$\bullet$ We theoretically investigate the robustness of distributed DNNs against adversarial action. We leverage notions of IB theory and rigorously prove that the robustness of distributed DNNs is affected by the splitting point and latent dimension. While existing work often optimize the splitting point and feature dimension to minimize end-to-end latency of distributed computing [7, 29], for the first time we reveal that they are also key factors of robustness which need to be considered in designs;
$\bullet$ We perform extensive experiments with the ImageNet-1K [30] dataset, by considering 6 different DNN architectures, 6 different distributed DNN approaches under 10 different attacks to support our theoretical findings. The results show that the theoretical analysis applies to the experimental settings under consideration. We share our code for reproducibility at https://github.com/asdfqwezxcf/AdvLatent, and we hope this work may open the door to a new field dedicated to studying the resilience of distributed DNNs.
This paper is organized as follows. Section 2 define the threat model under consideration. Next, Section 3 presents our theoretical analysis based on IB. Section 4 discusses our experimental setup while Section 5 presents our experimental results. Finally, Section 6 summarizes the related work and Section 7 draws conclusions and discusses possible directions for future work.
2 Threat Model
Figure 2: Threat model under consideration. (Left) The adversary plays a man-in-the-middle attack where the communication between mobile and local devices are altered without detection; (Right) Difference between adversarial attack in input space and adversarial attack in latent space.
Overview of the Adversarial Scenario. Figure 2 shows the threat model under consideration. We consider a man-in-the-middle attack to distributed DNNs where the mobile device and local server are connected through an insecure network. In this scenario, attackers can intercept transmissions and modify communications between mobile and edge endpoints without detection. More specifically, the attacker can add unperceptive perturbations to latent representations to mislead the local DNN at the edge, compromising the task-oriented performance of the distributed computing system. This attack differs fundamentally from traditional adversarial attacks in input space as latent representations often contains richer semantic information (e.g. objects and shapes) and less task-irrelevant information (e.g., background noise) [31], resulting in distinctive robustness characteristics compared to raw input data.
Analog to conventional adversarial attacks in the input space, we consider $l_{p}$ bounded attacks in our threat model. Remark that while $l_{p}$ constraints are first proposed to model the human invisibility in vision tasks [24], it has been widely applied in diverse tasks that operate without human involvement, such as cybersecurity [32] and wireless communication [33]. This is because the $l_{p}$ distortion represents a worst-case scenario for resilience. Thus, we believe $l_{p}$ bounded attacks is an essential starting point to assess the robustness of distributed DNNs.
Adversarial Attacks in Input Space. Let $f:\mathbb{R}^{d}\mapsto\mathbb{C}^{k}$ denote a DNN where $\mathbb{R}$ and $\mathbb{C}$ are respectively the input and output space, and $d$ and $k$ are the corresponding dimension of these two spaces. The DNN will assign highest score to the correct class $y=\operatorname*{arg\,max}_{k}f(x)$ for each input $x$. The adversarial goal is to introduce a perturbation $\delta_{d}\in\mathbb{R}^{d}$ to the original sample so that
$\operatorname*{arg\,max}_{k=1,...,K}f(x+\delta_{d})\neq y,$ (1)
where $||\delta_{d}||_{p}\leq\sigma$ and $\sigma$ is the distance constraint under different $l_{p}$ norm. Additionally, for visual applications, $\delta_{d}$ should satisfy the condition $x+\delta_{d}\in[0,1]^{d}$ as there is an explicit upper and lower bound for red, green and blue (RGB) value in digital images.
Adversarial Attacks in Latent Space. Let $g:\mathbb{R}^{d}\mapsto\mathbb{H}^{t}$ and $f:\mathbb{H}^{t}\mapsto\mathbb{C}^{k}$ denote the mobile DNN and local DNN, where $\mathbb{H}$ and $t$ are the latent space and its associated dimension, respectively. For each input $x$, the mobile DNN will generate a corresponding latent representation $g(x)\in\mathbb{H}^{t}$ and the local DNN will generate output $y=arg\max_{k}f(g(x))$ by taking the latent representation as input. Adversarial action in latent space adds a perturbation $\delta_{t}\in\mathbb{H}^{t}$ such that
$\operatorname*{arg\,max}_{k=1,...,K}f(g(x)+\delta_{t})\neq y,$ (2)
where $||\delta_{t}||_{p}\leq\sigma$ is the distance constraint under $l_{p}$ norm. We remark that the latent representations are model-dependent and there is no explicit bound for their value other than their computer-level representation (e.g., float, integer, double).
3 Theoretical Analysis
3.1 Analysis through Information Bottleneck
Figure 3: Modeling DNN with IB. Each representation $T_{i}$ only depends on the previous output $T_{i-1}$, and the optimal $T^{*}_{i}$ can be interpreted as the IB which optimizes Equation 3 at layer $i$.
The Information Bottleneck (IB)is a model-agnostic information-theoretical framework introduced by [13] to extract the relevant information about a random variable (r.v.) $Y$ from another r.v. $X$ by finding a representation $T$ which compresses the information of $X$ while captures only the sufficient information about $Y$. As shown in Figure 3, we model a DNN with a Markov chain $Y\mapsto X\mapsto T_{1}\mapsto\cdots\mapsto T_{k}\mapsto\hat{Y}$, where $X$, $Y$, $\hat{Y}$ and $T_{i}$ are respectively the input, its label, the inference output and the output of the $i$-th hidden layer. The IB optimizes the following:
$\min_{P(T_{i}|X)}I(X;T_{i})-\beta\cdot I(Y;T_{i}),1\leq i\leq k$ (3)
where $I(X;T_{i})$ is the mutual information between $X$ and $T_{i}$ while $I(Y;T_{i})$ is the mutual information between $Y$ and $T_{i}$. Each layer can be thus described by its own unique information plane $(I(X;T_{i}),I(Y;T_{i}))$ which represents its compression and generalization capability. Notice that optimizing Equation 3 is equivalent to minimizing $I(X;T_{i})$ – i.e., learning to compress – whereas maximizing $I(Y;T_{i})$ – i.e., learning to generalize. To simplify notation, and without loss of generality, henceforth we will consider a single generic hidden layer $T$.
Lemma 1 (Latent Robustness).
For a given DNN, the mutual information $I(Y;T)$ quantifies the robustness at the layer $T$.
Proof.
The end-to-end robustness of the DNN model can be modeled with $I(Y;\hat{Y})$, which measures the mutual information between the label $Y$ and the DNN inference $\hat{Y}$ [34]. We apply the Data Processing Inequality (DPI) to describe the information loss during processing [35]:
$I(Y;X)\geq I(Y;T)\geq I(Y;\hat{Y})$ (4)
In short, the generalization metric $I(Y;T)$ of hidden layer $T$ also describes the upper bound of $I(Y;\hat{Y})$, which is intrinsically a measure of robustness at layer $T$. ∎
By assuming that adversarial perturbations are not observable (i.e., adversarial perturbation doesn’t change the ground-truth of mutual information), it follows there is a prior yet unknown optimal solution $I^{*}(Y;T)$ for a specific DNN architecture that satisfies the IB where adversarial perturbations cannot decrease the performance – in other words, $I^{*}(Y;T)$ is resilient to adversarial attacks. The key issue is that although each DNN has a hypothesis set defined by its parameters, the optimum parameter set exhibiting the largest $I^{*}(Y;T)$ is unknown. To this end, each trained DNN using a finite dataset $(X,Y)$ has its own estimation $I(Y;T)$ and adversarial perturbations can decrease the estimated mutual information by introducing distribution shift in dataset.
Theorem 1 (Compression-Robustness Trade-Off).
The adversarial robustness $I(Y;T_{adv})$ is jointly determined by its ideal generalization capability $I^{*}(Y;T)$ and dimensionality $\mathcal{O}(|\mathcal{T}||\mathcal{Y}|/\sqrt{n})$:
$I(Y;T_{adv})=I^{*}(Y;T)-\mathcal{O}(\frac{|\mathcal{T}||\mathcal{Y}|}{\sqrt{n}}),$ (5)
where $n$ denotes the number of data and $|\mathcal{T}|$, $|\mathcal{Y}|$ are the cardinality of $T$ and $Y$, respectively.
Proof.
[36] have proven that the estimated mutual information $I(Y;T)$ using a dataset $(X,Y)$ with finite samples has the following error bound:
$||I^{*}(Y;T)-I(Y;T)||\leq\mathcal{O}(\frac{|\mathcal{T}||\mathcal{Y}|}{\sqrt{n}}).$ (6)
As the adversarial examples aim to decrease the performance, the adversarial robustness $I(Y;T_{adv})$ approaches the lower bound in Equation 6. Thus, Equation 5 holds. ∎
Theorem 1 is the information version of the complexity-generalization trade-off in Probably Approximately Correct (PAC) learning. A larger latent space $|\mathcal{T}|$ (i.e., a more complex hypothesis set in PAC learning) will have larger variance resulting in decreased performance with inputs coming from a distribution different than $X$. Conversely, with a smaller latent space (i.e., a smaller hypothesis set), the DNN has more bias, which leads to less accuracy (i.e., a smaller $I^{*}(Y;T)$).
Since the IB is model-agnostic and generalized, we point out that our theoretical framework alo provides value to explain the robustness in the general adversarial setting. For example, [37] derived a similar theoretical finding that robustness of DNNs decreased with growing data dimension. However, they solely concentrate on dimensionality while our analysis delves into the trade-off. [38] and [39] empirically explored the trade-off between robustness and performance, a concept close to Theorem 1. However, we offer a rigorous theoretical explanation of this trade-off, revealing that dimensionality serves as the determining factor.
In distributed DNNs, bottlenecks are often introduced to compress the latent dimension in order to reduce the communication overhead. Theorem 1 reveals that the feature compression technique can also affect the resilience of distributed DNN. Bottlenecks enhance the adversarial robustness by reducing the variance $\mathcal{O}({|\mathcal{T}||\mathcal{Y}|}/{\sqrt{n}})$ but also introduce vulnerability by unintentionally adding bias to decrease $I^{*}(Y;T)$. Therefore, we believe Theorem 1 provides a novel perspective to achieve robust distributed DNN.
Lemma 2 (Information Distortion).
In a Markov chain $Y\mapsto X\mapsto T$, the conditional mutual information $I(Y;X|T)$ quantifies the distortion incurred during the transmission of information from $X$ to $T$.
Proof.
Following [34], we use Kullback-Leibler (KL) divergence $D_{KL}[P(Y|X)||P(Y|T)]$ to characterize the distortion in IB framework. The expectation of $D_{KL}$ is
$\mathbb{E}\{D_{KL}\}=\sum_{X,T}P(X,T)\sum_{Y}P(Y|X)\log\frac{P(Y|X)}{P(Y|T)}.$ (7)
By applying the conditional independence of Markov chain, Equation 7 can be refined to
$\mathbb{E}\{D_{KL}\}$ $=\sum P(X,T)P(Y|X,T)\log\frac{P(Y|X)}{P(Y|T)}$ (8)
$=\sum P(X,T,Y)\log\frac{P(Y|X)P(X|T)}{P(Y|T)P(X|T)}$
$=\sum P(X,T,Y)\log\frac{P(X,Y|T)}{P(Y|T)P(X|T)}$
$=I(X;Y|T).$
Theorem 2 (Input vs Latent Robustness).
With same level of distortion introduced by input $X_{adv}$ and latent $T_{adv}$, it has
$I(Y;T^{\prime})\leq I(Y;T_{adv}),$ (9)
where $T^{\prime}$ denotes the latent corresponding to $X_{adv}$.
Proof.
Due to the chain rule of mutual information,
$I(X;Y|T)=I(X,T;Y)-I(Y;T).$ (10)
For a Markov chain $Y\mapsto X\mapsto T$, the joint distribution $P(X,Y,T)$ has following property
$P(X,Y,T)$ $=P(T|X,Y)P(Y|X)P(X)$ (11)
$=P(T|X)P(Y|X)P(X).$
Therefore, $I(X,T;Y)$ can be simplified as
$I(X,T;Y)$ $=\mathbb{E}\left\{\log\frac{P(X,T,Y)}{P(X,T)P(Y)}\right\}$ (12)
$=\mathbb{E}\left\{\log\frac{P(T|X)P(Y|X)P(X)}{P(T|X)P(X)P(Y)}\right\}$
$=\mathbb{E}\left\{\log\frac{P(Y|X)}{P(Y)}\right\}=I(X;Y).$
From Equation 10 and Equation 12, it follows that
$I(X;Y|T)=I(X;Y)-I(Y;T).$ (13)
Let $Y\mapsto X_{adv}\mapsto T^{\prime}$ be the Markov chain of adversarial attacks in input space, the distortion introduced by $X_{adv}$ is
$I(X_{adv};Y|T^{\prime})=I(X_{adv};Y)-I(Y;T^{\prime}),$ (14)
where $X_{adv}$ represents adversarial samples and $T^{\prime}$ is the corresponding latent representation.
Similarly, for a Markov chain $Y\mapsto X\mapsto T_{adv}$, we have
$I(X;Y|T_{adv})=I(X;Y)-I(Y;T_{adv}),$ (15)
where $T_{adv}$ and $I(X;Y|T_{adv})$ represents the latent perturbation and corresponding information distortion.
Assume $X_{adv}$ and $T_{adv}$ have same level of distortion,
$I(X_{adv};Y)-I(Y;T^{\prime})=I(X;Y)-I(Y;T_{adv}).$ (16)
Since $X_{adv}$ is a mapping of $X$, there is a Markov chain $Y\mapsto X\mapsto X_{adv}$. By DPI, it follows that
$I(X;Y)\geq I(X_{adv};Y).$ (17)
Therefore, Equation 9 holds. ∎
Theorem 2 formally describes that with same level of information distortion, attacking the latent space is less effective than attacking the input space. An important corollary can be derived from Theorem 2 by generalizing the input space to any latent space that is before $T$ in the Markov chain.
Corollary 1 (Depth-Robustness Trade-Off).
Assume same level of distortion is introduced to two latent space $T_{i-1}$ and $T_{i}$ where $i$ denotes the depth of DNN layers, it has
$I(Y;T_{i}^{\prime})\leq I(Y;T^{i}_{adv}),$ (18)
where $T_{i}^{\prime}$ denotes the latent at layer $i$ after introducing perturbation to $T_{i-1}$ while $T^{i}_{adv}$ denotes the latent at layer $i$ directly perturbed by distortion.
The proof is the same as Theorem 2 by replacing the Markov chain $Y\mapsto X\mapsto T$ to $Y\mapsto T_{i-1}\mapsto T_{i}$. Corollary 1 describes that the latent space in early layers will be more vulnerable than the the deeper latent space. This is because a small information loss gets amplified during propagation. We remark that this theoretical finding is in line with [40], which states that the difference between adversarial and benign samples is less significant in early layers compared to deeper ones, thus enabling the early exits.
Corollary 1 is critical to distributed DNNs as finding the optimal splitting point (depth of latent space) of DNN is one of the fundamental research questions to optimize the computation and communication overhead of distributed computing [7, 29]. However, we theoretically demonstrate that the splitting point is also a key factor of robustness which needs to be considered in design.
3.2 Connect Theory to Experiments
One of the open challenges in IB is to properly estimate the mutual information between high dimensional data. Existing literature usually involves simple DNN architecture and small datasets [39, 41]. To this end, our large-scale empirical study in Section 5 is based on end-to-end performance rather than IB. In order to connect the theoretical framework to experiments, we perform an experiment on MNIST with a simple DNN that consists of 3 layers of convolutional neural networks (CNNs) followed by a linear layer. Each CNN layer is followed by a batch normalization and a ReLU activation. Maxpooling layers are applied after the first and second CNN layer to downsample the features and a global average pooling layer is applied before the linear layer. The simple DNN achieved near perfect accuracy (99.42%) on MNIST.
Table 1: Estimated mutual information $\hat{I}(Y;T)$ and classification accuracy caused by input and latent perturbations as a Function of $\epsilon$ on MNIST.
MINE [42] NWJ [43] CPC [44] CLUB [45] DoE [46] Accuracy
Input Latent Input Latent Input Latent Input Latent Input Latent Input Latent
$\epsilon=0.01$ 2.01 2.08 1.28 1.31 2.25 2.26 9.88 9.88 27.44 27.88 98.68% 98.89%
$\epsilon=0.02$ 2.01 2.06 1.25 1.28 2.23 2.24 9.82 9.83 26.47 27.37 97.09% 97.89%
$\epsilon=0.03$ 1.93 1.95 1.17 1.26 2.19 2.22 9.74 9.77 25.28 26.88 94.27% 96.24%
$\epsilon=0.04$ 1.83 1.98 1.10 1.22 2.16 2.20 9.64 9.69 23.82 26.26 88.79% 93.40%
$\epsilon=0.05$ 1.78 1.88 1.01 1.16 2.11 2.16 9.51 9.59 22.03 25.61 78.64% 88.63%
$\epsilon=0.06$ 1.71 1.85 0.89 1.12 2.04 2.12 9.36 9.46 19.88 24.86 56.39% 79.96%
$\epsilon=0.07$ 1.62 1.80 0.79 1.12 1.94 2.07 9.19 9.32 17.13 24.06 31.59% 67.98%
$\epsilon=0.08$ 1.48 1.79 0.67 1.03 1.83 2.00 9.01 9.16 13.77 23.12 13.54% 52.26%
$\epsilon=0.09$ 1.35 1.65 0.56 1.00 1.69 1.91 8.81 8.98 9.83 22.22 4.34% 34.75%
$\epsilon=0.10$ 1.19 1.62 0.45 0.97 1.51 1.81 8.61 8.77 5.08 21.17 1.22% 20.58%
We consider the feature map output by the 2nd CNN layer as the target latent space and $l_{\infty}$ constrained Projected Gradient Descent (PGD) [17] with a perturbation strength $\epsilon$ is applied to both input and latent space. To have a comprehensive evaluation, multiple commonly used mutual information estimators are applied [42, 43, 44, 46, 45]. Among them, MINE [42], NWJ [43], and CPC [44] are lower bounded estimators while DoE [46] and CLUB [45] are upper bounded estimators.
Table 1 demonstrates the results of estimated mutual information $\hat{I}(Y;T)$ and corresponding classification accuracy as a function of noise level $\epsilon$. Note that the estimated mutual information varies significantly across different estimators, as they are built on different lower or upper bounds. While these estimations cannot give us exact value of mutual information, all of them show a same trend. To this end, we consider these estimations to be valid. As shown in Table 1, with an increasing perturbation budget $\epsilon$, all estimated $\hat{I}(Y;T)$ as well as the accuracy of CNN decrease. This indicates that the $I(Y;T)$ is an appropriate tool to quantify the DNN robustness as proposed in Lemma 1. Moreover, the mutual information as well as the classification accuracy associated with latent perturbations shows a higher value compared to those of input perturbations, which supports Theorem 2. This experiment bridges the theoretical analysis to large scale experiments in the rest of paper.
4 Experimental Setup
4.1 Attacks Under Consideration
Adversarial attacks can be categorized as gradient-based, score-based and decision-based approaches. Gradient-based attacks craft the adversarial samples by maximizing classification loss with gradient updates. On the other hand, score-based attacks leverage the output score of DNN as feedback to iteratively search the perturbation in the black-box setting and decision-based attacks can only access the hard label which has the highest score in the output. To have a thorough assessment, we implemented 10 popular attacks to DNNs. These include 4 gradient-based white-box attacks [14, 15, 16, 17], as well as 3 score-based black-box attacks [18, 19, 20] and 3 decision-based black-box attacks [21, 22, 23].
White-box Attacks: We consider 4 gradient-based attacks only in the white-box setting because latent representations are different for each model, resulting in numerous surrogate DNNs training in black-box scenarios that may be infeasible in practice. We choose Fast Gradient Sign Method (FGSM) [14], Basic Iterative Method (BIM) [15], Momentum Iterative Method (MIM) [16] with $l_{\infty}$ norm constraints. PGD [17] is implemented for both $l_{2}$ and $l_{\infty}$ spaces as a baseline for other black-box attacks.
Black-box Attacks: We consider 3 score-based attacks Natural Evolutionary Search (NES) [18], N-Attack [19] and Square Attack [20] in $l_{\infty}$ space and 3 decision-based attacks Evolutionary Attack (EVO) [21], Sign-OPT (S-OPT) [22] and Triangle Attack [23] in $l_{2}$ space.
Dataset and Metrics: We evaluate adversarial robustness using 1000 samples from the validation set of ImageNet-1K [30], limiting the samples to those which are correctly classified. We define the perturbation budget $\epsilon$ as the mean square error (MSE) under the $l_{2}$ norm constraint (i.e., $\epsilon\times d=\sigma^{2}$ and $\epsilon\times t=\sigma^{2}$ in input and latent space respectively where $d$, $t$ and $\sigma$ are the dimension of input, dimension of latent, and the distance constraint defined in Section 2.) and the maximum element-wise distance under $l_{\infty}$ norm constraint (i.e., $\epsilon=\sigma$), we define the attack success rate (ASR) as
$\mbox{ASR}(\epsilon)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{I}\left\{\operatorname*{arg\,max}_{k=1,\dots,K}f(x_{i},\delta_{i})\neq y_{i}\right\},$ (19)
where $\mathbf{I}\{\cdot\}$ is the indicator function and $f(x_{i},\delta_{i})$ is the DNN output when fed with the $i$-th sample.
Table 2: List of feature compression approaches considered in this paper
Category Approach                                  Description
Dimension Reduction SB naive supervised compression trained with cross entropy [8, 11]
DB bottleneck trained with naive knowledge distillation [9]
BF multi-stage training with distillation and cross entropy [7]
Datasize Reduction JC reduce precision in frequency domain using JPEG approach [47]
QT uniformly compress every element using naive bit quantization [27]
Advanced ES bottleneck trained with distillation and information-based loss
and data compressed with quantization and entropy coding [28]
4.2 DNNs Under Consideration
DNN Architectures. First, we consider 3 DNNs: VGG16 from [24] as well as ResNet50 and ResNet152 from [25]. Both VGG and ResNet have 5 feature extraction blocks which consist of several identical convolutional layers to extract features in different depths. During experiments, we consider the output of the third block as the target latent space without further specification. (We investigate the latent robustness as a function of depth in Section 5.4.)
Next, to investigate the effect introduced by the feature compression layer (i.e., the “bottleneck”) proposed for distributed DNNs, we introduce the same bottleneck design as [29] to VGG16, ResNet50 and ResNet152 and denote the new architectures as VGG16-fc, ResNet50-fc and ResNet152-fc.
Compression Approaches. In distributed DNNs, compression can be categorized as dimension reduction and data size reduction. Dimension reduction applies the bottleneck layer to compress the cardinality of the latent space while data size reduction leverages lossless (e.g. entropy coding) or lossy (e.g. quantization, JPEG) coding to minimize the bit rate of latent representations. Note that dimension reduction is a standalone approach that can applied on top of the data size reduction and vise versa.
We consider 3 different bottleneck training strategies for dimension reduction: Supervised Bottleneck (SB) [8, 11], Distillated Bottleneck (DB) [9] and BottleFit (BF) [7]. We also choose 2 data size reduction strategies JPEG Compression (JC) [47], Quantization (QT) [27] as well as 1 advanced approach Entropic Student (ES) that both compress the dimension and data size [28]. We summarize these approaches in Table 2.
5 Experimental Results
5.1 Performance w.r.t DNN Architecture
We first assess the latent robustness against different attack algorithms. Figure 4 shows the ASR obtained on ResNet152-fc with perturbation budget $\epsilon=0.01$. Remarkably, we notice that the ASR is higher for attacks in the input space than attacks in the latent space for each attack algorithm considered. In the case of Triangle Attack, the latent ASR is 88% less than the input ASR. On average, the ASR in input is 57.49% higher than the ASR obtained by attacks in the latent space. Moreover, Square Attack, EVO, and Triangle Attack have lowest ASR on latent representations. This is because these attacks search perturbations in a lower dimensional space, and hence it is more challenging for the adversary to find the effective distortions in compressed latent space. Figure 4 support the finding of Theorem 2, which states that under the same distortion constraint, the latent space constantly demonstrates better robustness than the input space.
Figure 4: 10 different attacks to ResNet152-fc with perturbation budget $\epsilon=0.01$.
As depicted in Figure 4, although various attacks achieved different ASR, they also demonstrated a similar trend in performance loss. For this reason, in the following experiments we only show results of selected attacks for briefness. To investigate the effect of different DNN architectures, we choose PGD as the white-box baseline, N-Attack as the score-based attack, and Triangle Attack as the representative of decision-based attack. Figure 5 shows the ASR of PGD, N-Attack and Triangle Attack on different DNNs with $\epsilon=0.003$. For each DNN, the ASR is higher in input-space attacks. In VGG16-bf, which shows the best robustness, the average ASR in the latent space is 87.8% lower than input attacks. On average, latent representations are 58.33% more robust. Therefore, the experiments in Figures 4 and 5 demonstrate that Theorem 2 is valid across various adversarial attacks and DNN architectures.
5.2 Performance w.r.t Compression Approach
Figure 5: Whitebox baseline (PGD) and blackbox attacks under $l_{\infty}$ and $l_{2}$ in input and latent space with perturbation budget $\epsilon=0.003$ applied to 6 different DNNs.
To evaluate the robustness of different compression approaches, we choose Square Attack and Triangle Attack which are more recent approaches for score-based and decision-based attacks respectively. We do not consider gradient-based attacks as compression approaches such as DB, QT, ES can lead to gradient masking. Hence, their robustness cannot be correctly evaluated by naive gradient-based attacks [48]. We choose a larger perturbation budget ($\epsilon=0.05$) than the experiments depicted in Figure 4 to further evaluate whether the compressed feature space is robust to attacks relying on low-dimensional sub-space searching. We note that data compression can be applied in addition to bottlenecks. However, for ablation study purposes, we choose ResNet50 without bottlenecks for JC and QT and ResNet50-fc for others.
Figure 6 shows the ASR of Square Attack in $l_{\infty}$ space and Triangle Attack in $l_{2}$ space with perturbation budget $\epsilon=0.05$. Except the JC and QT, the adversarial robustness shows the same trend regardless of the examined approaches. The average ASRs in input space are 79.07% and 87.22% higher than the average ASRs in latent space for Square Attack and Triangle Attack respectively. For DNNs with bottlenecks, despite the increase in perturbation, the ASR of Square Attack and Triangle Attack performed in latent representations do not increase distinctively comparing to Figure 4. However, since JC and QT do not have separate feature compression layers, the ASR of Square Attack and Triangle Attack in input space are only 55.8%, 0.25% higher than the attacks in latent space, showing a significant downgrade comparing to the other 4 approaches. These results confirm that the compressed feature space is indeed robust to attacks that search in lower dimensions.
Figure 6: Square and Triangle attack success rate associated with 6 feature compression approaches with perturbation budget $\epsilon=0.05$.
5.3 Performance w.r.t Dimension
In the previous experiment, we have shown that the robustness of the latent representation is mostly characterized by the bottleneck layer properties rather than the data size compression approach. Thus, we further evaluate the robustness for different sizes of latent space using N-Attack, MIM under $l_{\infty}$ constraint and S-OPT, PGD under $l_{2}$ constraint with multiple perturbation budgets ($\epsilon=0.003$; $\epsilon=0.01$; $\epsilon=0.03$). The cardinality of the latent space is controlled by the number of channels at the bottleneck layer. We first set the channel number as 12 for ResNet152-fc that can achieve 77.47% validation accuracy, which is almost similar to the performance of the original ResNet152 (78.31%). Then, we reduce the channel size to 3, which decreases the dimension of latent representations but also reduces the end-to-end performance to 70.01%. We do not repeat the results for Square Attack and Triangle Attack since they fail to achieve satisfactory ASR in the previous experiments due to their smaller search subspace, as shown in Figures 4 and 6.
Figure 7 shows results obtained by considering the $l_{\infty}$ and $l_{2}$ attacks with multiple perturbation budgets ($\epsilon=0.003$; $\epsilon=0.01$; $\epsilon=0.03$) in the latent space of original ResNet-152, 12-channel ResNet152-fc and 3-channel ResNet152-fc. From ResNet152 to 12-channel ResNet152-fc, the ASR reduces as the dimensionality of latent representations decreases. However, after reducing the channel size to 3, the ASR does not decrease any further. Conversely, distributed DNNs with a smaller channel size become more vulnerable to perturbations. (One exception is MIM with $\epsilon=0.03$, where the ASR of the 12-channel model is $79.5\%$ and the ASR of the 3-channel model is $75.0\%$. However, due to the difference of devices and random seeds, the ASR can vary 2-3%. Thus we do not consider the decrease of the MIM success rate in 3-channel ResNet152-fc which is less than 5%.) This is because when reducing channels from 12 to 3 channels, the accuracy also decreases to 7.46%, which in turn lessens the end-to-end generalization capability (i.e., $I^{*}(Y;T)$). This experiment supports our analysis in Theorem 1 that the robustness in latent representations of distributed DNN is jointly determined by the end-to-end performance and feature dimensions.
Figure 7: $l_{\infty}$ and $l_{2}$ attack success rate for different latent cardinalities of ResNet152-fc with different perturbation budgets: (left) $\epsilon=0.003$; (center) $\epsilon=0.01$; (right) $\epsilon=0.03$.
5.4 Performance w.r.t Depth
To validate the Corollary 1, we evaluate the robustness in latent representations with respects to different depths by showcasing a small model VGG16 and a deeper model ResNet152. In both VGG16 and ResNet152, 5 feature extraction blocks which consist of several identical convolutional layers are used to extract features in different depths. We choose the first, the third and the last feature map denoted as Feature 0, Feature 2 and Feature 4, respectively. We choose PGD and Square Attack as the white- and black-box representatives respectively. To have a more comprehensive assessment, we explore the performance with multiple perturbation budget $\epsilon$.
As shown in Table 3, latent representations are always more robust than the input irrespective of the depth and perturbation strength, which is in line with our experiments in Section 5.1. Moreover, ASR decreases with the depth of feature maps for both VGG16 and ResNet152. Therefore, this experiment empirically reveal the trade-off between robustness and computational burden which is formally discussed in Corollary 1. Splitting DNNs in early layers can result in less computation in the mobile device but the latent space is more vulnerable to adversarial perturbations.
Table 3: ASR as a function of depth
Input Feature 0 Feature 2 Feature 4
PGD Square PGD Square PGD Square PGD Square
$\epsilon=0.003$ 98.8% 48.2% 50.3% 3.4% 22.3% 3.4% 6.5% 0.6%
$\epsilon=0.006$ 99.3% 79.1% 81.6% 9.3% 50.1% 8.6% 12.4% 1.4%
VGG16 $\epsilon=0.009$ 99.8% 91.2% 93.7% 13.4% 71.5% 12.5% 19.4% 2.5%
$\epsilon=0.012$ 99.7% 95.4% 96.9% 17.3% 87.2% 16.0% 27.5% 3.7%
$\epsilon=0.015$ 99.9% 98.8% 98.5% 22.1% 93.5% 20.0% 35.2% 4.5%
$\epsilon=0.003$ 88.1% 36.4% 65.5% 14.7% 43.1% 3.7% 1.6% 0.0%
$\epsilon=0.006$ 96.9% 63.3% 90.5% 31.0% 76.7% 8.4% 3.4% 0.2%
ResNet152 $\epsilon=0.009$ 98.4% 81.9% 96.9% 45.7% 92.9% 13.4% 4.9% 0.3%
$\epsilon=0.012$ 99.0% 90.0% 99.0% 56.7% 97.4% 17.7% 6.0% 0.6%
$\epsilon=0.015$ 99.4% 94.7% 99.1% 68.8% 98.5% 21.9% 7.3% 0.7%
6 Related Work
Distributed Neural Networks. There is an increasing need for DNNs-based computation on mobile devices. Lightweight DNNs specifically tailored for mobile devices [4, 49, 50] fail to achieve comparable performance while edge computing approaches may incur in excessive latency [51]. As an intermediate option, [52] divide a DNN into two parts executed at the mobile device and edge, respectively. To reduce the networking load due to the large latent space of DNNs, literature introduces a “bottleneck” layer before the division point to reduce the dimension of latent representations [8, 11, 26]. Whereas, naive bottlenecks suffer from noticeable task-specific performance loss. As such, a key research problem in distributed DNN is to optimize the trade-off between the latent dimension and end-to-end performance. To achieve lower latent dimension while preserving accuracy, recent work utilizes more advanced training such as knowledge distillation [12] and multi-stage training [7]. Nevertheless, current work only focuses on the trade-off between the communication load and accuracy but ignore the robustness aspect.
In addition, different splitting points can affect the compression performance and computational burden in resource-constrained devices. For example, some studies [27, 53, 54] introduce bottlenecks in deeper layers and optimize a trainable rate-distortion metric. While such studies can achieve a remarkable compression performance, they neglect the limited computational capability of mobile devices which may result in larger delay for computation. In contrast, [28] introduce bottleneck in early layers with learnable prior for entropy coding to optimize the trade-off between compression and computation. [29] further investigate the effect of different splitting point in [28]. Our work reveals that the splitting point of DNN is highly related to the robustness to adversarial attacks in the distributed scenario.
Empirical and Theoretical Study of DNN Robustness. [55] propose a set of criteria to study adversarial robustness with empirical results, while [56, 57, 58] evaluate the robustness of different defense approaches. However, in the absence of a theoretical framework, different studies may arrive at conflicting conclusions. For example, [38] argue that there is a trade-off between generalization and robustness while [59] state that generalization does not affect the robustness. [39] argue adversarial training sacrifice accuracy on standard datasets while [60] can achieve same accuracy on both adversarial and standard datasets. On the other hand, PAC learning has been used to analyze the adversarial robustness of DNNs [61, 62, 63, 64, 65, 66]. However, such work mainly attempts to find minimum dataset size for a desired robustness level [61, 63, 66, 62]. Conversely, our investigation aims to evaluate adversarial robustness in DNNs that are trained with certain datasets. In addition, these approaches are evaluated on simple DNNs [64, 65], while we evaluate our findings on state-of-the-art DNNs. [34] first propose to use IB [13] to analyze DNNs and [31, 41] analyze the generalization and compression capability of DNNs with experiments. However, due to the challenge of mutual information estimation [67], such studies only perform experiments on relatively small neural networks. Recent work applying IB for robust learning [68, 69, 70, 71] has shown significant enhancements. However, it does not involve a general analysis for robustness of DNNs.
7 Concluding Remarks
This paper has investigated adversarial attacks to latent representations of DNNs. First, we have theoretically analyzed the robustness of latent representations based on the information bottleneck theory. To prove our theoretical findings, we have performed an extensive set of experiments with 6 different DNN architectures, 6 different distributed DNN approaches and considering 10 different attacks in literature. Our investigation concludes that latent representations in deeper layers are more robust than those in early layers assuming the same level of information distortion. Moreover, the adversarial robustness in latent space is jointly determined by the feature size and the end-to-end model generalization capability. We hope that this work will inspire future work on the topic of adversarial machine learning on distributed DNNs.